A spelunker is surveying a cave. He follows a passage that goes 178 m straight west, then 214 m in a direction 45
Answer #1

1st Displacement

D1=-178i

2nd Displacement

D2=214sin45*i-214cos45*j =151.32 i -151.32 j

3rd displacement

D3=283sin30*i +283cos30*j =141.5 i +245.1 j

4th displacement

D4=X i+ yj

The sum of 4 displacement should be zero

D1+D2+D3+D4=0

-178i +151.3 i-151.3 j +141.5 i+245.1 j+X i +y j=0

X i+ Yj=-114.8 i -93.8 j

a)

D4=sqrt[114.82+93.82] =148.25 m

b)

Direction =tan-1(-93.8/-114.8)=39.25o